OpenStudy (anonymous):

4x^2-5y^2-16x-30y-9=0 What is the standards form of the equation of the conic given above?

OpenStudy (anonymous):

help!(:

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

first, do you know the type of conic this equation represents?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

this equation is in general form. we need to get it to standard form by completing the square. group all terms with x's together and y's together with parenthesis. what you got?

OpenStudy (anonymous):

do we factor them to when group together?

OpenStudy (anonymous):

that's later... just group them according to what i said.... also, move the constant over to the right side of the equal sign...

OpenStudy (anonymous):

so we have.. (4x^2-16x)-(5y^2+30y)=-9

OpenStudy (anonymous):

my bad..positive 9

OpenStudy (anonymous):

ok.... now factor out a 4 from the x group then factor out a 5 from the y group. what u got now?

OpenStudy (anonymous):

4x(x^2-4)-5y(y-6)=9

OpenStudy (anonymous):

no... just a 4 and just a 5....

OpenStudy (anonymous):

oh so 4(x^2-4x)-5(y^2+6y)=-9

OpenStudy (anonymous):

yes, that's it... and btw, the right side should be positive 9, right?

OpenStudy (anonymous):

oh yes! idk why I put negative. so where do we go from here?

OpenStudy (anonymous):

ok... this is the completing the square part...

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

what should the value of c be in order for the polynomial to be a perfect square trinomial? \(\large x^2 -4x + c \)

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

HINT: take half of the x coefficient and square it. that's what c should be.

OpenStudy (anonymous):

yes... it's 4.. so...

OpenStudy (anonymous):

so is( x-4)^2

OpenStudy (anonymous):

\(\large x^2-4x +\color {red}4=(x-2)^2 \) agreed?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok... so back to the original problem...

OpenStudy (anonymous):

4(x-2)^2-5(y+3)^2=9

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

\(\large 4(x^2-4x+\color{red}4 )-5(y^2+6y)=9\color{red} {+ 16}\) since you added 4 to the x group to make it a perfect square, do you see why you needed to add 16 to the right side?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok... so when you do the same for the y group, what do you need to add to the right side also?

OpenStudy (anonymous):

-45

OpenStudy (anonymous):

yes... so what should the right side be then?

OpenStudy (anonymous):

-20

OpenStudy (anonymous):

yes... lemme summarize what you have so far...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\(\large 4(x^2-4x+4 )-5(y^2+6y+9)=9+ 16-45 \) \(\large 4(x-2)^2-5(y+3)^2=-20 \) now, standard form of a conic has a "1" on the right hand side... divide both sides by -20. what u got?

OpenStudy (anonymous):

(x-2)^2/5-(y+3)^2/4=1

OpenStudy (anonymous):

\(\large 4(x-2)^2-5(y+3)^2=-20 \) \(\large \frac{4(x-2)^2}{-20}-\frac{5(y+3)^2}{-20}=\frac{-20}{-20} \) simplify that last line...

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

um.. you y group should be positive, and the x group should be negative...

OpenStudy (anonymous):

ok..i just switched them

OpenStudy (anonymous):

can you help we with another one similar?

OpenStudy (anonymous):

\(\large \frac{\cancel{4}(x-2)^2}{\cancel{-20}-5}-\frac{\cancel5(y+3)^2}{\cancel{-20}-4}=1 \)

OpenStudy (anonymous):

cleaned up it should be: \(\large \frac{(y+3)^2}{4}-\frac{(x-2)^2}{5}=1 \) which is an equation of a hyperbola.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw... if you have a similar one, post it as a separate question and someone should help u out... :)